OpenStudy (friendzone):

Solve the equation for x. -6x + 8x = -46 A) -23 B) 3.2 C) 12 D) 16

OpenStudy (friendzone):

Daniel's basic cell phone rate each month is $29.95. Add to that $5.95 for voice mail and $2.95 for text messaging. This past month Daniel spent an additional C dollars on long distance. His total bill was $62.35. How much did Daniel spend on long distance? A) $23.50 B) $24.00 C) $62.35 D) $63.00

OpenStudy (friendzone):

f 5 + Y = 12, and you add -5 to the left side of equation, what should you add to right side of equation? A) -7 B) -5 C) 5 D) 7 Bonus Question 4) A store is having a sale on 200 DVD movies. Some of the movies are being sold for $10 and some of them are being sold for $12. If the store made $2200 after selling all 200 DVDs, how many $12 DVDs were sold? A) 50 B) 75 C) 100 D) 150

pooja195 (pooja195):

1) combine like terms and divide by the x value

OpenStudy (friendzone):

Solve. 3(x + 1) - 2x = -6. A) x = 1 B) x = 5 C) x = -7 D) x = -9 6) 3x + 2y = 2 -2x + y = 8 Solve the system of equations. A) (-2,4) B) (4,-2) C) (-4,-2) D) ( 5 4 , - 1 2 ) 7) Solve for x: 4x + 12 = 8 (given) 4x = -4 (subtraction) x = -1 What is the reason for the last step in the argument? A) division B) addition C) subtraction D) symmetric property 8) 2(x + 7) + 3x = 12 (given) 2x + 14 + 3x = 12 (?) 5x + 14 = 12 (simplify) 5x = -2 (subtraction) x = - 2 5 (division) What is the missing reason? A) Addition B) Subtraction Eliminate C) Reflexive Property D) Distributive Property

OpenStudy (friendzone):

2x + 5 = 10 What is the first operation used in solving this equation? A) addition B) division C) multiplication D) subtraction 10) Jeremy sells pencils. His daily sales have been linear. Jeremy sold 20 pencils on the first day, 23 pencils on the second day, and 26 pencils on the third day. A function can model Jeremy’s sequence. What is the slope of that function? A) -20 B) -3 Eliminate C) 0 D) 3

OpenStudy (anonymous):

Which one of your questions have you tried working out?

OpenStudy (anonymous):

No offense but I will not work these out until you tell me what you think the answer is. I have spent years learning all of the math I have. This isn't a Google service

OpenStudy (ybarrap):

For all of these problems, you need to isolate your variable by subtraction and then division. That's essentially the formula for most of these. Also, think of items as variables. Then apply my previous advice.